Eight readings were measured between two points. The sample mean distance was found to be 79.674m. Distance readings had a sample standard deviation of ±2.5 mm. What is the 99% confidence interval for the population mean? Provide the confidence interval with mm precision.
Answer #1

Sample Mean,    x̅ = 79.674m = 79674mm

Level of Significance ,    α =    0.01
sample std dev ,    s =    2.5
Sample Size ,   n =    8

degree of freedom=   DF=n-1=   7
't value='   tα/2=   3.4995
      
Standard Error , SE =   s/√n =   0.8839
margin of error ,   E=t*SE =   3.0931
confidence interval is       
Interval Lower Limit=   x̅ - E =    79670.9069 mm
Interval Upper Limit=   x̅ + E =    79677.0931 mm